OpenStudy (anonymous):

ok, so sorry, buuuuut..... I need help with this desperately... 2. Methane (CH4), ammonia (NH3), and oxygen (O2) can react to form hydrogen cyanide (HCN) and water according to this equation: CH4 + NH3 + O2  HCN + H2O You have 8 g of methane and 10 g of ammonia in excess oxygen. Answer the following questions: • What is the balanced equation for this reaction? • Which reagent is limiting? Explain why. • How many grams of hydrogen cyanide will be formed? Show your work.

OpenStudy (anonymous):

1) Balanced equation : 2CH4 + 2NH3 +302 --> 2HCN + 6H20 2) CH4=8/16=0.5 NH3=10/17=0.59 (No. of moles) One mole of CH4 requires one mole of NH3. 0.5 moles of CH4 requires only 0.5 moles of NH3. But there are 0.59 moles of NH3 hence NH3 is in excess and CH4 is limiting reagent. 3) 0.5*27= 13.5 g.

OpenStudy (anonymous):

thank you xD